OpenStudy (meilendurcer):

If a ping pong ball falls from the top of a building that is 168 feet high, how many seconds will it take the ball to hit the ground? Remember, the formula for distance is d = 16t^2.

OpenStudy (anonymous):

set \[168=16t^2\] and solve for \(t\)

OpenStudy (anonymous):

divide by 16 and get \[10.5=t^2\] and so \(t=\sqrt{10.5}\)

OpenStudy (meilendurcer):

good i had that answer. The volume of a box can be found by multiplying the length times width times height of a box (V = lwh). If the volume of a box is 1,300 cubic inches, and all its dimensions are equal, what is the side length of the box in inches?

OpenStudy (anonymous):

take the cubed root of 1300

OpenStudy (meilendurcer):

36?